### v2

@@ 2 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol and its security under unentangled attackers
    Each party performs a POVM $\{A^{x}_{a}\}$ and $\{B^{x}_{b}\}$, respectively, to the state $\rho$ and they send an answer $a,b\in\{0,1,\perp\}$ to their corresponding verifier.
The loss of quantum messages can be taken as an advantage from the point of view of the attackers. For example, assume that under the execution of the protocol the honest parties expect that half of the photons (or more) are lost in the transmission, i.e., $\eta\leq\frac{1}{2}$, but the attackers are able to receive all photons (since they can position themselves closer to the verifiers). Then the attacker Alice could make a random guess $\tilde{x}$ for $x$, measure the qubit $Q$ in the $\tilde{x}$ basis and broadcast the outcome and $\tilde{x}$ to fellow attacker Bob. After 1-round of simultaneous communication with Bob, they both know if the guess was correct. If so, they send the outcome to the verifiers, otherwise, they claim no photon arrived. Notice that Alice’s basis guess will be correct half of the time and therefore, the attackers can respond correctly half of the time, which is the expected ratio of transmission. We call this strategy the guessing strategy, and we denote it by $\mathbf{S}_{guess}$, see below.
@@
The tuple $\{\rho,A^{x}_{a},B^{x}_{b}\}_{x,a,b}$ will be called a non-entangled (NE) strategy. The probabilities that the verifiers, after the attackers’ actions (for the random variable $\textsc{V}_{\textsc{AB}}$) record correct, wrong, no photon and different answers (the verifiers ‘Abort’ in such a case) are thus, respectively, given by
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]$ $=\frac{1}{\lvert\mathcal{X}\rvert}\sum_{a\in\{0,1\},x\in\mathcal{X}}\mathrm{Tr}\left[\rho V^{x}_{a}\otimes A^{x}_{a}\otimes B^{x}_{a}\right],$ (2)
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]$ $=\frac{1}{\lvert\mathcal{X}\rvert}\sum_{a\in\{0,1\},x\in\mathcal{X}}\mathrm{Tr}\left[\rho V^{x}_{a}\otimes A^{x}_{1-a}\otimes B^{x}_{1-a}\right],$ (3)
@@
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{No photon}]$ $=\frac{1}{\lvert\mathcal{X}\rvert}\sum_{x\in\mathcal{X}}\mathrm{Tr}\left[\rho\mathbb{I}\otimes A^{x}_{\perp}\otimes B^{x}_{\perp}\right],$ (4)
@@
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}]$ $=\frac{1}{\lvert\mathcal{X}\rvert}\sum_{a\neq b\in\{0,1,\perp\},x\in\mathcal{X}}\mathrm{Tr}\left[\rho\mathbb{I}\otimes A^{x}_{a}\otimes B^{x}_{b}\right].$ (5)
Through this paper, $\mathbb{P}[*]$ will denote the probability of event *. Similarly, we will consider the random variable $\textsc{V}_{\textsc{P}}$ if the actions are performed by the prover. We say that an attack to a round of the protocol is successful if the following are fulfilled:
@@
$\begin{split}&\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]=\eta(1-p_{err})=:p_{\textsc{c}},\hskip 36.98866pt\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]=\eta p_{err}=:p_{\textsc{w}},\\ &\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{No photon}]=1-\eta=:p_{\perp},\hskip 48.36958pt\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}]=0=:p_{\lightning}.\end{split}$ (6)
This defines a probability distribution (denoted as a vector) $\textbf{p}_{hp}=(p_{\textsc{c}},p_{\perp},p_{\textsc{w}},p_{\lightning})$ that the honest prover would reproduce.
@@
A single round attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta,f}$  for $\eta=1$ can be identified with a so-called monogamy-of-entanglement (MoE) game, introduced by Tomamichel, Fehr, Kaniewski and Wehner in [TFKW13], formalized below in Definition 2.3 and generalized by Johnston, Mittal, Russo, and Watrous [JMRW16]. The authors of [TFKW13] showed that the optimal probability that the attackers are correct in one round of the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol is $81-\cos^{2}\pi/$ 8. Moreover, they show strong parallel repetition, i.e. if $\mathrm{QPV}_{\mathrm{BB84}}$ is executed $n$ times in parallel, the probability that the attackers are correct is at most $(1-\cos^{2}\pi/8)^{n}$.
Here we consider an extension of a MoE game, which we will call lossy MoE game, that will capture a round attack of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$.
@@
A lossy monogamy-of-entanglement game with parameter $\eta\in[0,1]$ consists of a finite dimensional Hilbert space $\mathcal{H}_{V}$, corresponding to party $V$, and a list of measurements $\{V_{v}^{x}\}_{v\in\mathcal{V}}$ on $\mathcal{H}_{V}$, indexed by $x\in\mathcal{X}$, where $\mathcal{V}$ and $\mathcal{X}$ are finite sets. Two collaborative parties, Alice and Bob, with associated Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, prepare an arbitrary quantum state $\rho_{VAB}$ and send $\rho_{V}$ to $V$, holding on $\rho_{A}$ and $\rho_{B}$, respectively. $V$ chooses $x\in\mathcal{X}$ uniformly at random and measures $\rho_{V}$ using $\{V_{v}^{x}\}$ to obtain the measurement outcome $v$. Then she announces $x$ to Alice and Bob. The collaborative parties make a guess of $v$ and they win the game if and only if both either guess $v$ correctly or their strategy, see below, is such that (8) is fulfilled.
A MoE game as in [TFKW13] is recovered for $\eta=1$. The interpretation of the lossy MoE game compared with the original MoE game is that the two collaborative parties have the ‘extra power’ of answering $\perp$’ which can be seen as them saying ‘I do not know’ with probability $1-\eta$.
Definition 2.4.
A strategy $\mathbf{S}_{MoE}^{\eta}$ for a lossy monogamy-of-entanglement game is a tuple
$\mathbf{S}^{\eta}_{MoE}=\{\rho_{VAB},A_{v}^{x},B_{v}^{x}\}_{v\in\mathcal{V}\cup\{\perp\},x\in\mathcal{X}},$ (7)
where $\rho_{VAB}$ is a density operator on $\mathcal{H}_{V}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and for all $x\in\mathcal{X}$, $\{A_{v}^{x}\}_{v\in\mathcal{V}\cup\{\perp\}}$ and $\{B_{v}^{x}\}_{v\in\mathcal{V}\cup\{\perp\}}$ are POVMs on $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, such that
@@
$\frac{1}{\lvert\mathcal{X}\rvert}\sum_{x\in\mathcal{X}}\mathrm{Tr}\left[\rho_{VAB}\mathbb{I}\otimes A^{x}_{\perp}\otimes B^{x}_{\perp}\right]=1-\eta.$ (8)
In [TFKW13] it is shown that any strategy can be purified in the sense that, enlarging the corresponding Hilbert spaces if necessary, $\rho_{VAB}=\left|\psi\rangle\langle\psi\right|$ is a pure state for $\left|\psi\right\rangle\in\mathcal{H}_{V}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{B}$ and $\{A_{v}^{x}\}$, $\{B_{v}^{x}\}$ are projective measurements. Therefore, unless it is explicitly specified, from now on we will consider any strategy $\mathbf{S}^{\eta}_{MoE}$, with $\mathbf{S}^{\eta}_{MoE}$ pure. A MoE game can be associated with an attack to the $\mathrm{QPV_{BB84}}$ protocol in the sense that having a strategy to break the protocol implies having a strategy for the MoE game [TFKW13, Section 5]. In a similar manner, it follows that having a $\mathbf{S}^{\eta}_{MoE}$ strategy implies having a NE strategy for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$. The idea is that in item 2. in the attack described as Attack 2.2, the attackers start with a tripartite state shared among the verifier, Alice and Bob and their task is to correctly guess the measurement outcome of the measurement which is performed on the verifier’ register. In the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol case, $V$ corresponds to the verifiers with associated Hilbert space $\mathcal{H}_{V}=\mathbb{C}^{2}$, with $\mathcal{X}=\{0,1\}$ and $\mathcal{V}=\{0,1\}$. The verifiers $V$ perform the collection of measurements
$\{V_{0}^{x},V_{1}^{x}\}_{x\in\{0,1\}},$ (9)
where $V_{0}^{x}=H^{x}\left|0\rangle\langle 0\right|H^{x}$ and $V_{1}^{x}=H^{x}\left|1\rangle\langle 1\right|H^{x}$, where $H$ is the Hadamard transformation, and the two collaborative parties, who correspond to the attackers, want to break the protocol by guessing the verifier’s outcome. The strategy $\mathbf{S}_{\textsc{tfkw}}=\{\left|\psi\rangle\langle\psi\right|,A_{a}^{x}=\delta_{a0},B_{a}^{x}=\delta_{a0}\}$, where $\left|\psi\right\rangle=\cos\frac{\pi}{8}\left|0\right\rangle+\sin\frac{\pi}{8}\left|1\right\rangle$, gives the optimal probability of winning the MoE game (see discussion below) and thus the optimal probability of being correct attacking the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol,
@@
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]=\frac{1}{\lvert\mathcal{X}\rvert}\sum_{a,x}\mathrm{Tr}\left[\left|\psi\rangle\langle\psi\right|V^{x}_{a}\otimes A^{x}_{a}\otimes B^{x}_{a}\right]=1-\cos^{2}\frac{\pi}{8}.$ (10)
This strategy also gives, using (6),
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}]=\cos^{2}\frac{\pi}{8},\hskip 28.45274pt\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}]=0.$ (11)
Comparing it with (6), and considering that $\eta=1$, the attackers could successfully attack one round of the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol if $p_{err}=\cos^{2}\frac{\pi}{8}$. Notice that if in a lossy MoE game the attackers actually ‘play’, meaning that they do not respond $\perp$’, we have
$\begin{split}&\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]+\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]\\ &+\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]=1.\end{split}$ (12)
In fact, since for QPV we impose $\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Abort}]=0$, the above expression reduces to
$\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]+\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Wrong}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]=1.$ (13)
We define the probability of winning, $p_{win}$ as the maximum probability of being correct conditioned on answering, i.e. $p_{win}:=\max\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]$. In fact, this corresponds to what in [TFKW13] is defined as the probability of winning the MoE game. Moreover, since for the BB84 MoE game it is always the case that $\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}$, we have that the optimal probability of winning the game is
$p_{win}=\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}\mid\textsc{V}_{\textsc{AB}}\neq\textsc{No photon}]=\mathbb{P}[\textsc{V}_{\textsc{AB}}=\textsc{Correct}]=1-\cos^{2}\frac{\pi}{8},$ (14)
which is obtained using $\mathbf{S}_{\textsc{tfkw}}$.
The security of the $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol relies on the probability that the attackers can win the lossy MoE game. The usage of Semidefinite Programming (SDP) techniques to bound non-local games was initiated by Cleve, Hoyer, Toner and Watrous [CHTW04] and Wehner [Weh06]. In order to apply these techniques, we take our security approach on the probability that the attackers can actually play the game without being caught, i.e., the probability that they can answer, $p_{ans}$,

### v3

@@ 2 The $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ protocol
    Alice intercepts the qubit register $P$, and applies an arbitrary quantum channel $\mathcal{E}_{PA_{0}\rightarrow AB}$ to it, and to $\sigma_{A_{0}}$. The subscript $PA_{0}\rightarrow AB$ indicates that the map has input and output registers $PA_{0}$ and $AB$, respectively. Let $\rho_{V\!AB}$ be the resulting state, that is,
    ${\mathbb{I}_{V}\otimes\mathcal{E}_{PA_{0}\rightarrow AB}(\left|\Phi^{+}\rangle\langle\Phi^{+}\right|_{VP}\otimes\sigma_{A_{0}})=\rho_{V\!AB}}.$ (6)
    Alice possesses registers $A$ and $B$, and $V_{0}$ holds $V$. On the other side, Bob intercepts $x$, and copies it.
  3. 3.
    Alice keeps register $A$ of $\rho$ and sends register $B$ to Bob. Bob keeps a copy of $x$ and sends a copy to Alice.
  4. 4.
@@
    After one round of simultaneous communication, each party performs a POVM $\{A^{x}_{a}\}_{a\in\{0,1,\perp\}}$ and $\{B^{x}_{b}\}_{b\in\{0,1,\perp\}}$, on registers $A$ and $B$ of the state $\rho$, and they send answers $a,b$, respectively, to their corresponding closest verifier.
See Fig. 2 for a schematic representation of a generic attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ in the No-PE model. The tuple $\mathbf{S}_{\eta}:=\{\rho,A^{x}_{a},B^{x}_{b}\}_{x,a,b}$ will be called a strategy for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$.
$V_{0}$||||||||timeAlice||||Bob||||$V_{1}$||||$A$$B$$x$$x$$\mathbb{I}_{V}\otimes\mathcal{E}(\left|\Phi^{+}\rangle\langle\Phi^{+}\right|_{VP}\otimes\sigma_{A_{0}})=\rho_{V\!AB}$$x$position$\{A^{x}_{a}\}_{a}$$\{B^{x}_{b}\}_{b}$$\rho_{V\!AB}$
Figure 2: Schematic representation of a generic attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$ in the No-PE model. Straight lines represent classical information, and undulated lines represent quantum information. The gray-shaded region represents the tripartite quantum state $\rho_{VAB}$.
To illustrate how photon loss can be exploited by adversaries, consider the following example. Suppose the verifiers expect that an honest prover would have a transmission rate of at most $\eta\leq\frac{1}{2}$, meaning that half or more of the photons are lost in transit. In contrast, since the adversaries may position themselves closer to the verifiers and intercept most of the transmitted photons, we give them the extra power of intercepting all of them. The attack proceeds as follows: Alice selects a random bit $\tilde{x}\in\{0,1\}$, measures the received qubit in the $\tilde{x}$ basis, and forwards the outcome and $\tilde{x}$ to Bob. After a single round of simultaneous communication, both attackers know whether the guess was correct. If $\tilde{x}=x$, they return the measurement outcome; otherwise, they claim photon loss, i.e.  output $\perp$. Since the basis guess is correct with probability $1/2$, this strategy succeeds with probability $1/2$—matching the expected success rate of an honest prover under $\eta=\frac{1}{2}$. We refer to this strategy as the guessing strategy, denoted $\mathbf{S}_{\mathrm{guess}}$.
@@
The probabilities that the verifiers, after the attackers’ actions (for the random variable $\textsc{V}_{\textsc{AB}}$) record ‘Correct’, ‘Wrong, ‘No photon’, and ‘Abort’ (different answers) are thus, respectively, given by
$q_{\textsc{c}}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]$ $=\frac{1}{2}\sum_{a,x\in\{0,1\}}\mathrm{Tr}\left[\rho V^{x}_{a}\otimes A^{x}_{a}\otimes B^{x}_{a}\right],$ (7)
$q_{\textsc{w}}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\right]$ $=\frac{1}{2}\sum_{a,x\in\{0,1\}}\mathrm{Tr}\left[\rho V^{x}_{a}\otimes A^{x}_{1-a}\otimes B^{x}_{1-a}\right],$ (8)
@@
$q_{\perp}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\perp\right]$ $=\frac{1}{2}\sum_{x\in\{0,1\}}\mathrm{Tr}\left[\rho\mathbb{I}_{V}\otimes A^{x}_{\perp}\otimes B^{x}_{\perp}\right],$ (9)
@@
$q_{\lightning}:=\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\lightning\right]$ $=\frac{1}{2}\sum_{a\neq b\in\{0,1,\perp\},x\in\{0,1\}}\mathrm{Tr}\left[\rho\mathbb{I}\otimes A^{x}_{a}\otimes B^{x}_{b}\right].$ (10)
In every round $i\in\{1,\ldots,r\}$ of the protocol (executed sequentially $r$ times), the attackers will pick a strategy $\mathbf{S}^{i}_{\eta}$ that can depend on the previous rounds. A strategy $\mathbf{S}^{i}_{\eta}$ will induce a probability vector $\textbf{q}^{i}=(q_{\textsc{c}}^{i},q_{\perp}^{i},q_{\textsc{w}}^{i},q_{\lightning}^{i})$. In an honest implementation,
@@
$\begin{split}&\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]=\eta(1-p_{err})=:p_{\textsc{c}},\hskip 28.45274pt\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{w}\right]=\eta p_{err}=:p_{\textsc{w}},\\ &\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\perp\right]=1-\eta=:p_{\perp},\hskip 56.9055pt\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\lightning\right]=0=:p_{\lightning}.\end{split}$ (11)
This defines a probability vector $\textbf{p}_{hp}=(p_{\textsc{c}},p_{\perp},p_{\textsc{w}},p_{\lightning})$ that the honest prover would ideally reproduce. An attack is successful if the verifiers cannot distinguish if their data came from the distribution $\textbf{p}_{hp}...\textbf{p}_{hp}$ ($r$ times) or from $\textbf{q}^{1}...\textbf{q}^{r}$. As mentioned above, in Section 2.2 we provide a test to distinguish between these two cases based on the received data.
2.1 Exact loss-tolerance of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$
@@
A single round attack to $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$  for $\eta=1$ can be identified with a so-called monogamy-of-entanglement (MoE) game, introduced by Tomamichel, Fehr, Kaniewski and Wehner in [TFKW13], formalized below in Definition 2.5 and generalized by Johnston, Mittal, Russo, and Watrous [JMRW16]. The authors of [TFKW13] showed that the optimal probability that the attackers are correct in one round of the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol is $\cos^{2}(\frac{\pi}{8})$. Moreover, they show strong parallel repetition, i.e.  if $\mathrm{QPV}_{\mathrm{BB84}}$ is executed $m$ times in parallel, the optimal probability that the attackers are answer correct is $(\cos^{2}(\frac{\pi}{8}))^{m}$. Here we consider an extension of a MoE game, which we will call lossy MoE game, that will capture a round attack of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$—specifically step 4—and extensions of it, see Section 4.
@@
Let $\mathcal{X}$ and $\mathcal{V}$ be finite non-empty alphabets. Let $V^{x}_{v}$ be POVMs of the same finite dimension for all ${(x,v)\in\mathcal{X}\times\mathcal{V}}$, and let :=$\mathcal{M}:=\{V^{x}_{v}\}_{x,v}$. A lossy monogamy-of-entanglement game with parameter ${\eta\in[0,1]}$, played by a referee, with associated Hilbert space $\mathcal{H}_{R}$, and two collaborative parties Alice and Bob, denoted by
$\textsf{G}_{\eta}:=(\eta,\mathcal{M}),$ (12)
is described as follows:
  1. 1.
    Alice and Bob, with associated Hilbert spaces $\mathcal{H}_{A}$ and $\mathcal{H}_{B}$, respectively, prepare a quantum state $\rho_{RAB}\in\mathcal{S}(\mathcal{H}_{R}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{B})$.
  2. 2.
    They send register $R$ to the referee, holding on registers $A$ and $B$, respectively. The two parties are no longer allowed to communicate.
  3. 3.
    The referee chooses $x\in\mathcal{X}$ uniformly at random and measures register $R$ using $\{V_{v}^{x}\}_{v}$ to obtain the measurement outcome $v$. Then, he announces $x$ to Alice and Bob.
  4. 4.
    The collaborative parties make a guess for $v$ and they win the game if and only if both either guess $v$ correctly or both answer $\perp$ (with probability $1-\eta$). In order to obtain the answers, Alice and Bob perform POVMs $\{A_{a}^{x}\}_{a\in\mathcal{V}\cup\{\perp\}}$ and $\{B_{a}^{x}\}_{a\in\mathcal{V}\cup\{\perp\}}$ on their local registers, respectively. The tuple $\mathbf{S}_{\eta}:=\{\rho_{RAB},A_{v}^{x},B_{v}^{x}\}_{v\in\mathcal{V}\cup\{\perp\},x\in\mathcal{X}}$ will be called a strategy for $\mathcal{G}_{\eta}$.
The lossy constraint, i.e.  answering $\perp$ with probability $1-\eta$, is given by
@@
$\frac{1}{\lvert\mathcal{X}\rvert}\sum_{x\in\mathcal{X}}\mathrm{Tr}\left[\rho_{RAB}\mathbb{I}_{V}\otimes A^{x}_{\perp}\otimes B^{x}_{\perp}\right]=1-\eta.$ (13)
See Fig. 3 for a schematic representation of a lossy MoE game. A monogamy-of-entanglement game $\mathsf{G}$ as introduced in [TFKW13] is recovered by setting $\eta=1$, i.e.  $\mathsf{G}=\mathsf{G}_{\eta=1}$. The winning probability of a lossy MoE game $\mathsf{G}_{\eta}$, given a strategy ${\mathbf{S}_{\eta}=\{\rho_{RAB},A_{v}^{x},B_{v}^{x}\}_{v\in\mathcal{V}\cup\{\perp\},x\in\mathcal{X}}}$, is given by
@@
$\omega(\mathsf{G}_{\eta},\mathbf{S}_{\eta})=\frac{1}{\lvert\mathcal{Z}\rvert}\sum_{v,x}\mathrm{Tr}\left[\rho_{RAB}V^{x}_{v}\otimes A^{x}_{v}\otimes B^{x}_{v}\right].$ (14)
The optimal winning probability is given by the supremum of (14) over all possible strategies, i.e.
$\omega(\mathsf{G}_{\eta}):=\sup_{\mathbf{S}_{\eta}}\omega(\mathsf{G}_{\eta},\mathbf{S}_{\eta}).$ (15)
$x$$x$$x$$R$$\{V_{v}^{x}\}_{v}$$v$$A$$\{A_{a}^{x}\}_{a}$$a$$B$$\{B_{b}^{x}\}_{b}$$b$$\rho_{RAB}$
Figure 3: Schematic representation of a lossy monogamy-of-entanglement game. The gray-shaded region represents the tripartite quantum state $\rho_{RAB}$ prepared by Alice and Bob and shared amongst the three parties. The referee, Alice and Bob are denoted by $R$, $A$ and $B$, respectively.
As shown in [TFKW13], any strategy can be purified in the sense that, by enlarging the Hilbert spaces if necessary, one may assume $\rho_{RAB}=\left|\psi\rangle\langle\psi\right|_{RAB}$ for some pure state $\left|\psi\right\rangle\in\mathcal{H}_{R}\otimes\mathcal{H}_{A}\otimes\mathcal{H}_{B}$, and that the local measurements $\{A_{a}^{x}\}_{a}$ and $\{B_{b}^{x}\}_{b}$ are projective for all $x\in\mathcal{X}$. From now on, we will assume our strategies are of this purified form.
Next, we describe the $\mathsf{G}^{BB84}$ game, originally introduced in  [TFKW13], which was used by the authors to show security of $\mathrm{QPV}_{\mathrm{BB84}}$ in the No-PE model. We will later use its lossy version to show security of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$.
Example 2.6.
The BB84 monogamy-of-entanglement game is described as follows. Alice and Bob prepare a quantum state and send a qubit from it to the referee, who chooses uniformly at random to measure the qubit either in the computational or the Hadamard basis. Upon knowing the choice of basis, the task of Alice and Bob is to guess the measurement outcome. Using the above terminology, the game is given by
$\mathsf{G}^{BB84}=\left(\eta=1,\mathcal{M}\right),$ (16)
where =$\mathcal{M}=\{V_{0}^{x},V_{1}^{x}\}_{x\in\{0,1\}},$ with
$V_{0}^{x}=H^{x}\left|0\rangle\langle 0\right|H^{x}\text{ and }V_{1}^{x}=H^{x}\left|1\rangle\langle 1\right|H^{x},$ (17)
where $H$ is the Hadamard transformation. Varying $\eta\in[0,1]$ defines the lossy BB84 MoE game, denoted by $\mathsf{G}^{BB84}_{\eta}$.
The $\mathsf{G}^{BB84}$ game can be associated with an attack to the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol in the sense that having a strategy to break the protocol in the No-PE model implies having a strategy for the MoE game [TFKW13, Section 5], and therefore
$\mathrm{Pr}\left[\textsc{V}_{\textsc{AB}}=\textsc{c}\right]\leq\omega(\mathsf{G}^{BB84}).$ (18)
In a similar manner, it follows that having a strategy $\mathbf{S}^{BB84}_{\eta}$ for $\mathsf{G}^{BB84}_{\eta}$ implies having a No-PE strategy for $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$. The idea is that in step 4 in the attack described above, the attackers start with a tripartite state shared among the verifier, Alice and Bob and their task is to correctly guess the measurement outcome of the measurement which is performed on the verifier’s register. In the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol case, verifier $V_{0}$ plays the role of the referee with associated Hilbert space $\mathcal{H}_{V}=\mathbb{C}^{2}$, with $\mathcal{X}=\{0,1\}$ and $\mathcal{V}=\{0,1\}$. In the purified version of $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, $V_{0}$ performs the measurement =$\mathcal{M}=\{V_{0}^{x}=H^{x}\left|0\rangle\langle 0\right|H^{x},V_{1}^{x}=H^{x}\left|1\rangle\langle 1\right|H^{x}\}_{x\in\{0,1\}}$, and the two collaborative parties, who correspond to the attackers, want to break the protocol by guessing the verifier’s outcome. We refer the reader to step 4 of the attack on $\mathrm{QPV}_{\mathrm{BB84}}^{\eta}$, illustrated in Fig. 2 and note that it corresponds to a lossy MoE game [^4] Although not represented in the figure, $V_{0}$ performs the measurement $\{V^{x}_{v}\}_{v}$ in his local register $V$ of the state $\rho_{VAB}$., depicted in Fig. 3. The strategy $\mathbf{S}_{\textsc{tfkw}}=\{\left|\psi\rangle\langle\psi\right|,A_{a}^{x}=\delta_{a0},B_{a}^{x}=\delta_{a0}\}$, where $\left|\psi\right\rangle_{V}=\cos\frac{\pi}{8}\left|0\right\rangle_{V}+\sin\frac{\pi}{8}\left|1\right\rangle_{V}$, gives the optimal probability of winning the $\mathsf{G}^{BB84}$ game [TFKW13] (see discussion below) and thus the optimal probability of being correct attacking the $\mathrm{QPV}_{\mathrm{BB84}}$ protocol is upper bounded by